a boat can travel 35 mph in still water. if it travels 117 miles with the current in the same length of time it travels 93 miles against the current, whats the speed of the current?
Answer by mananth(16949)

Boat speed = 35 mph
current speed = x mph

speed against current 35 - x
speed with current 35 + x

Distance against current 93 miles
Distance with current 117 miles

t=d/r time against = time with current -
93 / ( 35 - x )= 117 / ( 35 + x )
93 ( 35 + x )= 117 ( 35 -x )
3255 + 93 x = 4095 - -117 x
117 x 93 x 4095 + -3255
210 x = 840
/ 210
x = 4 mph current speed